OpenStudy (caitlinnr14):

–x + 6y = 30 Change into slope intercept

OpenStudy (anonymous):

-x + 6y = 30 6y = x + 30 6y/6 = x/6 + 30/6 can you finish it?

OpenStudy (caitlinnr14):

5 but where would you put that

OpenStudy (anonymous):

You need to change it into slope intercept. Slope intercept looks like y =mx + b

OpenStudy (caitlinnr14):

so the 5 would go in 6y=5x+30??

OpenStudy (anonymous):

So y = mx + b -x + 6y = 30 6y = x + 30 6y/6 = x/6 + 30/6 \[ y = \frac{1}{6}x + 5 \] Do you see?

OpenStudy (anonymous):

The answer to your question is \[y = \frac{1}{6}x + 5 \]

OpenStudy (caitlinnr14):

oh yes thx

OpenStudy (anonymous):

Yw